OpenStudy (anonymous):

Medal A patient with high blood pressure who weighs 230 pounds is put on a diet to lose 30 pounds in 3 months. The patient loses 8 3/4 pounds the first month and 11 1/2 pounds the second month. How much weight must be lost the third month for the goal to be achieved? ? lb

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

I re checked my answer the correct answer is 14 lb and 1/4

OpenStudy (anonymous):

I got this by subtracting 11.5 and 8.75 together and got 2.75 so i added 2.75 to the 11.5

OpenStudy (anonymous):

add the two together then subtract from 30 that is what is left to lose

OpenStudy (anonymous):

what two numbes do you add up

OpenStudy (anonymous):

8 3/4 and 11 1/2 you can change to decimals or change fractions to lcd

OpenStudy (anonymous):

the lowest common denominator would be 4 so it will be 11 2/4 + 8 3/4=?

OpenStudy (doc.brown):

\[230-8\frac{3}{4}-11\frac{1}{2}-x=230-30\]

OpenStudy (doc.brown):

\[230-230+30-8\frac{3}{4}-11\frac{1}{2}=x\\x=30-8\frac{3}{4}-11\frac{1}{2}\\x=\frac{4}{4}(30)-\left(\frac{4}{4}(8)+\frac{3}{4}\right)-\left(\frac{4}{4}(11)+\frac{2}{2}(\frac{1}{2})\right)\\x=\frac{120}{4}-\frac{35}{4}-\frac{46}{4}\\x=\frac{39}{4}\\x=9\frac{3}{4}\]

OpenStudy (anonymous):

thank you so much :)